In order to prepare pH = 7.00 of 1.00 L buffer solution, you added 12.43 g tris (MW: 121.135), calculate the mass in g of tris hydrochloride (MW: 157.596) that you will need to add. pKa = 8.072. Show your calculation process.
Answer #1

mass of tris = 12.43 g

moles of tris = 12.43 / 121.135 = 0.1026 mol

pH = 7.00

pKa = 8.072

pH = pKa + log [B / BH+]

7.00 = 8.072 + log [0.1026 / BH+]

moles of BH+ = 1.211 mol

mass of BH+ = 190.8 g

mass in g of tris hydrochloride = 190.8 g